From GlueXWiki
Jump to: navigation, search

July 17, 2014 FDC+CDC meeting


  1. DAQ (Beni)
  2. CDC status File:MStaibCDCUpdate17July2014.pdf (Mike, Beni)
  3. FDC status (Lubomir, Luke)
  4. Electronics (Chris, Nick)
  5. Engineering Cathode Scan system (Bill)
  6. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Curtis, Naomi (CMU), Sean (NU), Luke, Dave, Chris, Mike, Simon, Beni, Fernando, and Lubomir (JLab).

CDC update

Mike made a report, linked above:

- first slide shows the positions of the channels with noise as analyzed from the fADC data.

- next several slides demonstrate the "missing bit" problem. In one case all the channels in one module have that. Naomi: such problem was fixed by re-floating the circuit board.

- last slide shows the mean and the RMS of the pedestals. Questions about the meaning of the overflow/underflow in these histograms: Mike will work on that.

- Mike: the strain relief of the signal cables is not good enough: sometimes may lose contact. Fernando will find a way to improve it.

- According to Mike the pedestal subtraction in the Dave's code is done correctly, so this is not an explanation for the low CDC efficiency as discussed yesterday in the calibration meeting.

FDC update

- Lubomir presented a table (xls file linked) with all the problematic cathode channels as identified from the latest DAQ runs. Luke checked these channels with a scope which information is included in the Q-column with color. Information about Anatoly's measurements with scope in March/April (right after insertion) and the attempts to fix these channels is also included.

- For 69 of the channels (pink or red color) we suspect the problem is at the detector side (mostly conductive tape, but also connector or pre-amp issues). In the packages (1 to 4) the number of such channels is 4,1,41,23. Ten of these channels were OK in April, but had problems before, except two of them.

- We have 11 channels (green) where we see signals with the scope but not with fADC. However, the fADC channels are OK with another cable; Fernando suspect cable/connector problem. In the fADC data you see overflow or a fixed value at ~2000 and Beni made a list of such channels before.

- Finally, there are 9 channels (yellow) with no signal, but also almost no noise. Some of these were identified before to have problems with the 3M connectors, but the rest may have other issues.


- Fernando looked for EMI noise in the hall at the platform. He didn't find anything in the kHz range. The only source of very low noise was the mass flow meters for the BCAL Nitrogen. Dave proposed to check this by turning them off for a while. However, Fernando was not able to get very close to the CDC due to the target work there.

- Starting middle of August - September Fernando's group will start checking/fixing/re-routing cables that were not installed properly. Expect short interruptions at that time.


- Luke, with help from other people, started working on the scanning project (picture attached): the microscope has been attached to the motion stage, the motor can now be controlled from the computer. The shaft that connects the motor is missing but Tom ordered a new one; the encoder still needs to be connected.


- Beni:it's time to start using alcohol in the gas system. Will ask Scott to fill the alcohol bubblers.